OpenStudy (anonymous):

A ball is thrown vertically upward from the ground at a velocity of 150 feet per second. Its distance from the ground after t seconds is given by How fast is the ball moving 10 seconds after being thrown?

OpenStudy (anonymous):

so we know that \[v=v_{0}+at\]

OpenStudy (anonymous):

yes velocity

OpenStudy (anonymous):

or since this involves gravity our acceleration is constant so we have\[v=v _{0}-g t\]

OpenStudy (anonymous):

velocity = increment - gravity * time?

OpenStudy (anonymous):

so you would plug in the givens of 10s and initial velocity of 150ft/s which converts to about 45m/s

OpenStudy (anonymous):

and you have v=45-10*10

OpenStudy (anonymous):

velocity equals velocity initial times gravity time time

OpenStudy (anonymous):

times*

OpenStudy (anonymous):

oh so it is velocities starting value

OpenStudy (anonymous):

v0

OpenStudy (amistre64):

you are missing the "is given by" rule .... which then would have its derivative taken

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

Its distance from the ground after t seconds is given by _______

OpenStudy (anonymous):

ft

OpenStudy (anonymous):

given in feet lol

OpenStudy (anonymous):

toadill he is talking about how the question in the beginning says is given by and then it goes right to the next sentence of "How fast is..."

OpenStudy (anonymous):

How fast is the ball moving 10 seconds after being thrown

OpenStudy (amistre64):

recons physics formula is apt :)

OpenStudy (anonymous):

apt?

OpenStudy (anonymous):

yeah more errors

OpenStudy (amistre64):

apt: appropriate or suitable in the circumstances.

OpenStudy (anonymous):

I am just trying to review this is what the instructor gave me lol

OpenStudy (amistre64):

\[h(t) = -\frac12gt^2+v_ot+h_o\] \[h'(t) = v(t)\] \[v(t)=-gt+v_o\]

OpenStudy (anonymous):

by s(t) = -16t^2 + 150ft

OpenStudy (anonymous):

I guess it was an image instead of text.

OpenStudy (amistre64):

yeah, those copy pastes get eaten sometimes if they cant be processed correctly :)

OpenStudy (amistre64):

s(t) = -16t^2 + 150t , the derivative of displacement is velocity

OpenStudy (anonymous):

So you just find the derivative of the function?

OpenStudy (amistre64):

yes, and plug in t=10 for the solution

OpenStudy (anonymous):

okay thanks I am going to take a break now.

OpenStudy (amistre64):

margarita ville?

OpenStudy (anonymous):

yeah lol

OpenStudy (amistre64):

:) enjoy